the girl takes 2h to deliver newspapers to her 125 customers. the other girl delivers the same number of newspaper in 3h. how long will it take them to deliver the 125 newspapers if they work together?

Answer by checkley77(12569)

1/2+1/3=1/x
(3+2)/2*3=1/x
5/6=1/x cross multiply.
5x=6
x=6/5
x=1.2 hours working together.